OpenStudy (anonymous):

Find the vertex and the x intercept of the graph. y = x^2 + 2x + 1 I found the vertex, but I need help with the x-intercepts. The vertex is (-1,0). Using the formula x = -b/2a to find the vertex, plug in the numbers to get 2/2(1) which is -2/2 = -1. That is the x coordinate of the vertex. To find the y coordinate, you simply have to plug the x value into the original equation -1^2 + 2(-1)+ 1 is 1 - 2 + 1 = 0. Therefore, the vertex equals (-1, 0).

OpenStudy (anonymous):

x-intercept is always in the form of (x, 0), so y=0 for an x-intercept. x^2+2x+1=0 (x+1)^2=0 x=-1

OpenStudy (anonymous):

so, the x intercept (-1,0)

OpenStudy (anonymous):

yes, the x-intercept is also the vertex in this case.

OpenStudy (anonymous):

okay, thank you! :)

OpenStudy (anonymous):

it is 6x im sure